Ratio of rotational K.E. to rolling K.E. of a solid sphere is__________ ?

Grade:12

1 Answers

Vikas TU
3 years ago
If there is no sliding/slipping, then ω = v/r, where r is the radius of the sphere. The ratio of the rotational kinetic energy to the linear kinetic energy of a non-slipping uniform solid sphere is thus: (m*v²/5)/(m*v²/2) = 2/5